OpenStudy (anonymous):

Can someone help? (x+7)(x-1)=(x+1)^2 (x+7)(x-1)=(x+1)(x+1) 2x+6=2x+2 Is this break down correct so far?

OpenStudy (anonymous):

Now, how did you get (x+7)(x-1)=(x+1)^2?

OpenStudy (anonymous):

Or was that the original question?

OpenStudy (anonymous):

It is the original equation I am suppose to solve it

OpenStudy (anonymous):

Well the correct answer is 2

OpenStudy (asnaseer):

your breakdown is wrong

OpenStudy (anonymous):

Asnaseer were am I messing up at?

OpenStudy (asnaseer):

for example: (x+1)(x+1) = x*(x+1) + 1*(x+1) = x^2 + x + x + 1 = x^2 + 2x + 1

OpenStudy (asnaseer):

try applying the same rules to: (x+7)(x-1)

OpenStudy (anonymous):

My book shows the answer... but we are suppose to show the work.

OpenStudy (anonymous):

well that is 2x+6 right?

OpenStudy (asnaseer):

no - look at the way I expanded (x+1)(x+1) above

OpenStudy (asnaseer):

(x+7)(x-1) = x*(x-1) + 7*(x-1) try completing it from there

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

I am trying to see where you are coming up with that

OpenStudy (asnaseer):

ok, lets try it with actual numbers to help you understand this better...

OpenStudy (asnaseer):

8*5 = 40 so we can write this as: (1+7)(6-1) = 40 agreed?

OpenStudy (anonymous):

Agree

OpenStudy (asnaseer):

ok, now lets try to expand this: (1+7)(6-1) = 1*(6-1) + 7*(6-1) = 1*5 + 7*5 = 5 + 35 = 40 does that help?

OpenStudy (anonymous):

I am lost with this section = 1*(6-1) + 7*(6-1)

OpenStudy (anonymous):

So are we suppose to only multiply by the numbers in the first set of parenthesis

OpenStudy (anonymous):

If so then it makes sense but I didn't think that was how it was done.

OpenStudy (asnaseer):

ok, lets try another method. suppose we had: x + 7x do you agree this can be written as: (1 + 7)*x

OpenStudy (anonymous):

I don't see how that would be the same

OpenStudy (asnaseer):

well, now if we replace the x with (6-1) we get: (6-1) + 7(6-1) = (1+7)(6-1)

OpenStudy (asnaseer):

in general: (a+b)(c+d) = a(c+d) + b(c+d) = ac + ad + bc + bd

OpenStudy (anonymous):

okay..

OpenStudy (asnaseer):

if you are still stuck then try looking at this site: http://www.algebrahelp.com/lessons/simplifying/foilmethod/ it might help you understand better

OpenStudy (asnaseer):

do you think you can now expand this correctly? (x+7)(x-1) =

OpenStudy (anonymous):

Okay so (x+7)(x-1)=(x+1)^2

OpenStudy (anonymous):

=(x+1)^2 is 2x + 2 right?

OpenStudy (asnaseer):

no :(

OpenStudy (anonymous):

is it 2x+1

OpenStudy (asnaseer):

I would advise you to take a look at that site above first to get a better understanding

OpenStudy (anonymous):

I'm sorry I'm being so dumb. My class hasn't started yet but I have some health issues and I'm trying to get ahead start and my professor went ahead and sent me my assignments so I'm trying to learn this without having the class... =/

OpenStudy (asnaseer):

don't worry - you are certainly not dumb :) but if you really want to learn this I would spend some time on that site above, then come back here and we can continue.

OpenStudy (anonymous):

okay so the first part of the equation should be x^2-1x+7x+-7 ???

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

so then what would I need to do to do (x+1)^2

OpenStudy (asnaseer):

(x+1)^2 = (x+1)(x+1)

OpenStudy (anonymous):

x^2+x+x+1

OpenStudy (asnaseer):

perfect again! now you should first be able to simplify both expressions.

OpenStudy (asnaseer):

you know that +- turns into a minus correct?

OpenStudy (asnaseer):

and you can collect "like" terms

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so, for example: x^2-1x+7x+-7 = x^2 - x + 7x - 7 = x^2 + 6x - 7

OpenStudy (anonymous):

so that breaks down to x^2+6x-7=x^2+2x+1

OpenStudy (asnaseer):

thats right so far

OpenStudy (asnaseer):

now bring all the terms involving x from the right hand side to the left hand side - remember to change their signs when you do this

OpenStudy (anonymous):

x^2+4x=x^2+8

OpenStudy (asnaseer):

also bring over the x^2 term from the RHS

OpenStudy (anonymous):

I'm not sure how to handle x^2???

OpenStudy (anonymous):

4x=8 x=2

OpenStudy (asnaseer):

bingo! see you CAN do it! :)

OpenStudy (anonymous):

okay so now what would you do if only one side had x^2

OpenStudy (anonymous):

how would you break that down

OpenStudy (asnaseer):

before tackling that - one thing to remember is that you can check your answer by substituting x=2 into your original expressions

OpenStudy (anonymous):

yes.. I learned that with my other equations

OpenStudy (asnaseer):

good :)

OpenStudy (asnaseer):

now, if you did end up with some x^2 terms, then you would have what is known as a quadratic equation. have you studied these yet?

OpenStudy (asnaseer):

if not, this site should help you understand: http://www.mathsisfun.com/algebra/quadratic-equation.html

OpenStudy (anonymous):

wonderful! Thanks so much

OpenStudy (anonymous):

are you on here a lot?

OpenStudy (asnaseer):

there is also another very good site that is filled with lots of short video tutorials on various subjects here: http://www.khanacademy.org/

OpenStudy (asnaseer):

I am on this site quite a lot - but as I work full time, its usually in the evenings (UK time)

OpenStudy (anonymous):

Well would you mind if I needed help which I know I will if I messaged you?

OpenStudy (asnaseer):

there are many many helpful people on this site. I would suggest you just post your questions as usual and, only if you find no one is helping, then by all means send me a message and I'll do what I can to help out. it's great to find students like yourself that are eager to learn - you should be very proud of yourself :)

OpenStudy (anonymous):

Well thank you.. I get discouraged sometimes... I don't seem to get a lot of support from my family and often wonder if it will all pay off

OpenStudy (asnaseer):

have faith in your abilities - you are certainly a quick learner!

OpenStudy (anonymous):

if x=0 is that a no solution problem?

OpenStudy (asnaseer):

no - x=0 could well be a valid solution to some equation. e.g. x=0 is a solution to: 2x = 3x

OpenStudy (anonymous):

x(2x-3)=(2x+1)(x-4) 2x-3x=2x-8x+1x-4 -1x=-5x-4 4x=-4 x=-1

OpenStudy (anonymous):

Were am I messing up with this equation?

OpenStudy (anonymous):

where*

OpenStudy (asnaseer):

you made a mistake up there - remember that x * 2x = 2x^2

OpenStudy (asnaseer):

I need to go now - but good luck with the rest. remember to post each new question separately in the list to the left <----

OpenStudy (anonymous):

thanks so much!

OpenStudy (asnaseer):

yw :)